A body of mass 4kg moving with a velocity of 7m/s is acted upon by a force of 20N for 0.1s. Determine the final velocity of the body


1
Expert's answer
2021-11-24T19:10:51-0500

Solution.

"m=4 kg;"

"v_0=7 m\/s;"

"F=20N;"

"t=0.1 s;"

"Ft=mv-mv_0\implies v=\dfrac{Ft+mv_0}{m};"

"v=\dfrac{ 20N\sdot0.1s+4 kg\sdot 7 m/s}{4 kg}=7.5 m/s;"

Answer: "v=7.5 m\/s."